# Training-Conditional Coverage Bounds under Covariate Shift

### v3

@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
$\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\operatorname{\mathbb{P}}\left(n_{B}\geq n^{\prime}\right)\overset{(*)}{\geq}\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\left(1-2e^{-2n^{2\epsilon}}\right)$
@@
where ()(*)( ∗ ) holds according to Hoeffding’s inequality.
@@
and the inequalty 24 follows from the fact that $W^{<}(x)\leq W^{<}$ and $W^{>}(x)\leq W^{>}$ for all $x\in\mathbb{R}$. Hence,
@@
$1-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)},$
where in the last step we have used (the proof of) Durrett [2019][Theorem 2.2.11] with $\epsilon=\delta/2$ and $b_{n}=n$ twice; once for the first term with $X_{n,k}=\frac{dQ}{dP}(Z_{k})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{k})>B\right\}$, $k\in\mathcal{I}$, and the other time for the third term with $X_{n,k}=dQ/dP(Z_{k})$, $k\in\mathcal{I}$, and the requirements of the theorem are satisfied by $\operatorname{\mathbb{E}}dQ/dP(Z_{\text{calib}})=1<\infty$ according to Durrett [2019][Theorem 2.2.14]; also recall that $\mathcal{I}\subset[n]$ for the split conformal method corresponds to the calibration dataset. Putting everything together, we have
@@
$-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)}.$
@@ Appendix B Proof of Theorem 1
Recall from Section 2.3
@@
[…]
@@ B.1 Proof of Remark 1
[…]
@@ Appendix C Proof for Jackknife+
By McDiarmid’s inequality McDiarmid et al. [1989] we get
@@
$\ \leq 2p\exp\left(-\frac{2\kappa_{1}^{2}}{n}\left(\frac{\epsilon}{\kappa_{2}c_{n-1}}-\frac{1}{\kappa_{1}}\right)^{2}\right).$
@@ C.1 Proof of Theorem 3
$\supseteq\Bigg{\{}y\in\mathbb{R}:1-F^{(n-1)}\left(\left|y-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)>\alpha+\delta+2\epsilon L_{n-1}\Bigg{\}}$
@@
for ${\bf D}_{n}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}:=\left\{D:\max_{i}\|\mu_{\hat{\beta}_{-i}}-\mu_{\overline{\beta}_{-1}}\|_{\infty}<\epsilon\right\}$ and :=$\mathcal{B}:=\left\{D:\left\|\hat{F}^{(n-1)}-F^{(n-1)}\right\|_{\infty}<\delta\right\}$. From Lemma 6, we know
@@ C.2 Proof of Theorem 4
$\hat{\mathcal{C}}_{\alpha}^{\text{J+}}(X_{\text{test}})$ =$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}{w}_{i}\boldsymbol{1}\left\{\mu_{\hat{\beta}_{-i}}(X_{\text{test}})+\left|Y_{i}-\mu_{\hat{\beta}_{-i}}(X_{i})\right|<y\right\}<1-\alpha\right\}\bigcap$
@@
=$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}{w}_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{-i}}(X_{i})\right|<\left|y-\mu_{\hat{\beta}_{-i}}(X_{\text{test}})\right|\right\}<1-\alpha\right\}$
@@
=$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}\hat{w}_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{-i}}(X_{i})\right|\geq\left|y-\mu_{\hat{\beta}_{-i}}(X_{\text{test}})\right|\right\}>\alpha\right\}$
@@
$\supseteq\Bigg{\{}y\in\mathbb{R}:1-\hat{F}^{(n-1)}_{Q}\left(\left|y-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|+2\epsilon\right)>\alpha\Bigg{\}}.$
@@
$F_{Q}^{(n-1)}(t):=\operatorname{\mathbb{P}}_{Z_{1}\sim Q}\left(\left|Y_{1}-\mu_{\overline{\beta}_{-1}}(X_{1})\right|\leq t\right),$
@@
$\supseteq\Bigg{\{}y\in\mathbb{R}:1-F_{Q}^{(n-1)}\left(\left|y-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)>\alpha+\delta+2\epsilon L_{Q,n-1}\Bigg{\}}$
@@ D.1 Proof of Theorem 5
for ${\bf D}_{n}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}:=\left\{D:\|\mu_{\hat{\beta}_{n}}-\mu_{\overline{\beta}_{n}}\|_{\infty}<\epsilon\right\}$ and :=$\mathcal{B}:=\left\{D:\left\|\hat{F}^{(n)}-F^{(n)}\right\|_{\infty}<\delta\right\}$. From Lemma 6, we know $\operatorname{\mathbb{P}}({\bf D}_{n}\notin\mathcal{A})\leq 2p\exp\left({-\frac{2\kappa_{1}^{2}\epsilon^{2}}{n\kappa_{2}^{2}c_{n}^{2}}}\right)$. Also, according to Dvoretzky–Kiefer–Wolfowitz inequality, we have $\operatorname{\mathbb{P}}({\bf D}_{n}\notin\mathcal{B})\leq 2e^{-2n\delta^{2}}$. Thus,
@@ D.2 Proof of Theorem 6
$\hat{\mathcal{C}}_{\alpha}^{\text{full}}(X_{\text{test}})$ =$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}w_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{i})\right|<\left|y-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{\text{test}})\right|\right\}<1-\alpha\right\}$
@@
=$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}\hat{w}_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{i})\right|\geq\left|y-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{\text{test}})\right|\right\}>\alpha\right\}$
@@
$\operatorname{\mathbb{P}}\left({\bf D}_{n}\notin\mathcal{B}\left(\delta+2C\sqrt{\frac{B}{n}}\right)\right)\leq 4e^{-n\delta^{2}/(2B^{2})}$

### v4

@@ A.2 Weighted DKW Inequalities
$\geq\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\operatorname{\mathbb{P}}\left(n_{B}\geq n^{\prime}\right)$
$\overset{(*)}{\geq}\left(1-4e^{-n^{\prime}\delta^{2}/(2B^{2})}\right)\left(1-2e^{-2n^{2\epsilon}}\right)$
@@
where ()(*) holds according to Hoeffding’s inequality.
@@
and the inequalty 26 follows from the fact that $W^{<}(x)\leq W^{<}$ and $W^{>}(x)\leq W^{>}$ for all $x\in\mathbb{R}$. Hence,
@@
$1-2\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>n\right)-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}(Z_{\text{calib}})\right)^{2}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)},$ (27)
where in the last step we have used (the proof of) Theorem 9 with $\epsilon=\delta/2$ and $b_{n}=n$ twice; once for the first term with $X_{n,k}=\frac{dQ}{dP}(Z_{k})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{k})>B\right\}$, $k\in\mathcal{I}$, and the other time for the third term with $X_{n,k}=dQ/dP(Z_{k})$, $k\in\mathcal{I}$. The requirements of the theorem are satisfied by
$x\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{calib}})>x\right)\leq\operatorname{\mathbb{E}}\left(\frac{dQ}{dP}(Z_{\text{calib}})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})>x\right\}\right)=\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>x\right)\to 0$ (28)
as $x\to\infty$ and then according to Theorem 10 ; also recall that $\mathcal{I}\subset[n]$ for the split conformal method corresponds to the calibration dataset. Putting everything together, we have
@@
$-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}(Z_{\text{calib}})\right)^{2}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)}.$
@@ Appendix B Proof of Theorem 1
Recall from Section 2.3, we have
@@
where $Z_{i}\overset{\text{ i.i.d. }}{\sim}P$, $Z_{\text{test}}\sim Q$, and $\mathcal{I}$ denotes the set of indices corresponding to the calibration dataset. We note
@@ B.1 Proof of Remark 1
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\frac{2K(2C+1)}{\delta\sqrt{m}}\right)\leq\delta,\qquad\ 0<\delta\leq 1.$
@@ Appendix C Proof for Jackknife+
By McDiarmid’s inequality McDiarmid and others [1989] we get
@@
$\ \leq 2p\exp\left(-\frac{2\kappa_{1}^{2}}{n}\left(\frac{\epsilon}{\kappa_{2}c_{n-1}}-\frac{1}{\kappa_{1}}\right)^{2}\right),$
@@ C.1 Proof of Theorem 3
$\supseteq\Bigg\{y\in\mathbb{R}:1-F^{(n-1)}\left(\left|y-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)>\alpha+\delta+2\epsilon L_{n-1}\Bigg\}.$
@@
for ${\bf D}_{n}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}:=\left\{D:\max_{i}\|\mu_{\hat{\beta}_{-i}}-\mu_{\overline{\beta}_{-1}}\|_{\infty}<\epsilon\right\}$ and $\mathcal{B}:=\left\{D:\left\|\hat{F}^{(n-1)}-F^{(n-1)}\right\|_{\infty}<\delta\right\}$. From Lemma 6, we know
@@ C.2 Proof of Theorem 4
$\hat{\mathcal{C}}_{\alpha}^{\text{J+}}(X_{\text{test}})$ $=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}{w}_{i}\boldsymbol{1}\left\{\mu_{\hat{\beta}_{-i}}(X_{\text{test}})+\left|Y_{i}-\mu_{\hat{\beta}_{-i}}(X_{i})\right|<y\right\}<1-\alpha\right\}\bigcap$
@@
$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}{w}_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{-i}}(X_{i})\right|<\left|y-\mu_{\hat{\beta}_{-i}}(X_{\text{test}})\right|\right\}<1-\alpha\right\}$
@@
$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}\hat{w}_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{-i}}(X_{i})\right|\geq\left|y-\mu_{\hat{\beta}_{-i}}(X_{\text{test}})\right|\right\}>\alpha\right\}$
@@
$\supseteq\Bigg\{y\in\mathbb{R}:1-\hat{F}^{(n-1)}_{Q}\left(\left|y-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|+2\epsilon\right)>\alpha\Bigg\},$
@@
$\supseteq\Bigg\{y\in\mathbb{R}:1-F_{Q}^{(n-1)}\left(\left|y-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)>\alpha+\delta+2\epsilon L_{Q,n-1}\Bigg\}.$
@@ D.1 Proof of Theorem 5
for ${\bf D}_{n}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}:=\left\{D:\|\mu_{\hat{\beta}_{n}}-\mu_{\overline{\beta}_{n}}\|_{\infty}<\epsilon\right\}$ and $\mathcal{B}:=\left\{D:\left\|\hat{F}^{(n)}-F^{(n)}\right\|_{\infty}<\delta\right\}$. From Lemma 6, we know $\operatorname{\mathbb{P}}({\bf D}_{n}\notin\mathcal{A})\leq 2p\exp\left({-\frac{2\kappa_{1}^{2}\epsilon^{2}}{n\kappa_{2}^{2}c_{n}^{2}}}\right)$. Also, according to Dvoretzky–Kiefer–Wolfowitz inequality, we have $\operatorname{\mathbb{P}}({\bf D}_{n}\notin\mathcal{B})\leq 2e^{-2n\delta^{2}}$. Thus,
@@ D.2 Proof of Theorem 6
$\hat{\mathcal{C}}_{\alpha}^{\text{full}}(X_{\text{test}})$ $=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}w_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{i})\right|<\left|y-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{\text{test}})\right|\right\}<1-\alpha\right\}$
@@
$=\left\{y\in\mathbb{R}:\sum_{i=1}^{n}\hat{w}_{i}\boldsymbol{1}\left\{\left|Y_{i}-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{i})\right|\geq\left|y-\mu_{\hat{\beta}_{X_{\text{test}},y}}(X_{\text{test}})\right|\right\}>\alpha\right\}$
@@
Appendix E Experiment Details for Figure 1
The sizes of the training, calibration, and test datasets are 525, 225, and 247, respectively, for the Wine Quality, Abalone, and Combined Cycle Power Plant datasets. For the Concrete Compressive Strength dataset, the corresponding sizes are 360, 155, and 169. Distribution shift is introduced following the approach of Tibshirani et al. [2019], by resampling data points with probabilities proportional to $\exp(x^{\top}\beta)$, where
$\beta_{\text{wine quality}}=[0.5,1,0,0,0,0,0,0,0,0,0],$
$\beta_{\text{abalone}}=[0,10,10,0,0,0,0,0],$
$\beta_{\text{CCS}}=[0.01,0.01,0,0,0,0,0,0],$
$\beta_{\text{CCPP}}=[0.2,0,0,0].$
As a simple illustration of the role of likelihood ratio estimation, as discussed in Section 4, Figure 2 shows the impact of likelihood ratio estimation on the distribution of the training-conditional error for the Concrete Compressive Strength dataset.
Figure 2: The likelihood ratio is estimated using kernel density estimator with Gaussian kernel.